OpenStudy (anonymous):

What are the exact solutions of x2 = 5x + 2?

OpenStudy (anonymous):

\[x ^{2}-5x-2=0\] \[(-5)^{2}\pm \sqrt{(-5)^{2}-4(1)(-2)}/2\] \[(25\pm \sqrt{25+8})/2\] \[(25\pm \sqrt{33})/2\] So, the 2 exact solutions would be, \[(25+\sqrt{33})/2; (25-\sqrt{33})/2\] Alternatively, you can take square root of 33, so: \[(25+ 5.74)/2 ; (25-5.74)/2\] \[30.74/2 ; 19.26/2\] Final solutions are: 15.37 and 9.63